OpenStudy (anonymous):

Solve for x: −3(x + 3) = −3(x + 1) − 5

OpenStudy (anonymous):

you solution is not a solution, it doesnt solve the equation. Where is the value for x jhonyy?

jhonyy9 (jhonyy9):

-3x-9=-3x-3-5 -3x-9=-3x-8 add to both sides +3x -3x+3x-9=-3x+3x-8 -9=-8

jhonyy9 (jhonyy9):

the term of x is eliminated i think that you have missed anything

OpenStudy (anonymous):

-9=-8 doesnt make any sense, that is why there isnt a solution for this equation.

myininaya (myininaya):

no solution

myininaya (myininaya):

-9 does not equal -8